OpenStudy (anonymous):

What is the value of the expression 5 – c – d when c = 8 and d = –13? A. –26 B. –16 C. 0 D. 10

OpenStudy (alexandervonhumboldt2):

5-8-(-13)=-3+13=10 -(-)=+

OpenStudy (anonymous):

i got 100